SX FLAO Reconstructors with small bad actuator list using RR

04:30 Greg startup SX AO system

05:15 New AO Supervisor
  • Close loop but no slopes arriving at Switch BCU

05:30 Switch back to Old AO Supervisor
  • Close loop but not slopes

05:50 AO Cabinet is very warm
  • Switch BCU Stratix = 49.5

06:00 Steve tip over and open SX AO cabinet door
  • Switch temperature lower to 46

06:15 Old Supervisor
  • Close loop and slopes arrive

06:30 New Supervisor
  • No slopes
  • Restart adsec and wfs
  • no slopes
  • back to Old supervisor: slopes
  • back to New Supervisor: no slopes

07:00 try to put Old bad actuator list in new Supervisor
  • no slopes

07:20 Find that the fiber switch configuration had changes
  • Old Supervisor had the change in adsecarb.conf
  • New Supervisor was a copy of old supervisor before the change.
  • Copied old adsecarb.conf to new adsecarb.conf
  • Slopes now arriving

07:30 Use flat 20160218_flat3.sav
  • Close Loop
    • bin 1
    • 990 Hz
    • 400 modes
  • Success
  • Pupil look good with a small bit at 9:00
  • save flat 20170706_400_1.sav

07:45 Use new flat and close bin 1 loop
  • new flat
  • close loop in bin 2
  • new flat

08:00 Try to save Opt Loop Data with IRTC frames
  • When clicked save the WFS msg died
  • had to w_stop, w_start

08:17 Back up and running
  • Try to save IRTC data
  • "Saving Opt Loop Data failed" box popped up
  • IRTC did not save any frames.
  • Tried a second time

Collect Bin 1 Reconstructor

08:24 10 modes, Open Loop
  • gain file 600_optimized_6l
  • Disturb 082303
  • Intmat 082436
  • Recon 082611 (works in CL)

08:31 10 modes, CL
  • Disturb 082303
  • Intmat 083246
  • Recon 083415 (works in CL)

08:36 36 modes, CL
  • Disturb 083644
  • Intmat 083740
  • Recon 083919 (works in CL)

08:41 66 modes, CL
  • Disturb 084345
  • Intmat 084448
  • Recon 084632 (works in CL)

08:49 153 mode, CL
  • Disturb 084957
  • Intmat 085120
  • Recon 085244 (works in CL)

08:53 In Closed loop we see high currents on actuators 340 and 261. May have to put time in no ICC list.
  • These are where there are bits out of the pupil at 6:30 and 8:00

08:56 400 mode, CL
  • Disturb 085600
  • Intmat 085725
  • Recon 085900 (does not work in CL, edge rolls at 8:30)

09:03 CL with 153 mode recon 085244
  • flat 20170706_9
  • apply

09:07 Cl with 153 mode
  • see high current at 261 and 151
  • in flat have high current at 546

09:22 Set bad actuator list to IF list plus one found by Roberto to be bad (476)

09:31 CL
  • Tried 153 mode, got a big bite
  • Tried old 400 mode: Is OK
  • flat 20170706_400_10.sav
  • apply
  • Old 400 mode
  • flat 20170706_400_11.sav
  • apply high currents in (+) 546, 7 (-) 6

09:52 Set bad actuator list to FF plus one found by Roberto to be bad in pos (476) plus one causing high currents to in pos (6)
  • Still see high (+) currents at 564, 7 and high (-) currents at 228

09:58 400 mode CL
  • Disturb 085600
  • Intmat 095811
  • Recon 100118 (work in CL)
  • flat 20170706_400_14
  • apply

10:05 CL with new 400 mode
  • 200 Hz
  • 6 l/D
  • flat 20170706_400_15

10:11 CL with new 400 mode
  • 990 Hz
  • flat 20170706_400_16

10:16 add 7 and 37 to no_icc list
  • new 400 mode

10:26 add 564, 150 and 262 to no_icc list
  • new 400
  • flat _17

10:32 400 mode CL 400 mode 100118
  • Disturb 085600
  • Intmat 103146
  • Recon 103319
  • flat _18
  • a bit high currents 340, 261, 228, 149

10:44 CL new 400
  • flat _19

10:45 400 mode CL 400 mode 103319
  • Disturb 085600
  • Intmat 104657
  • Recon 104853
  • flat _20
  • apply

11:10 400 mode CL 400 mode old recon
  • Disturb 085600
  • Intmat 110855
  • Recon 111034
  • flat _22
  • cl
  • flat _24

11:16 Add 148 and 151 to no_icc list
  • apply _24
  • CL with 1112034
  • flat _25

11:32 400 mode with bigger disturbance
  • modeAmp 600_optimized_6l_1.5 (1.5 times larger then _6l)
  • Disturb 113238
  • Intmat 113339
  • Recon 113544
  • flat _26
  • apply
  • flat _27
  • 200 Hz, 6l/D
  • flat _28

11:54 990 Hz with latest Recon
  • flat _29
  • apply
  • flat

11:32 400 mode
  • 48 iterations
  • modeAmp 600_optimized_6l
  • Disturb 115845
  • Intmat 120215
  • Recon 121750

12:22 CL with new Recon
  • 1.2" seeing
  • 990 Hz
  • Opt gain failed

12:27CL with new Recon
  • 1.2" seeing
  • 990 Hz
  • 122729, 122808, 122853

  • 122808 Modal:
    122808_modal.jpg

12:27CL with new Recon
  • no seeing
  • 990 Hz
  • 123309, 123419, 123439

  • 123439 Modal:
    123439_modal.jpg

-- %USERSIG{DougMiller - 2017-07-06}%
I Attachment Action Size Date Who Comment
122808_modal.jpgjpg 122808_modal.jpg manage 139 K 06 Jul 2017 - 12:35 DougMiller 122808 Modal
123439_modal.jpgjpg 123439_modal.jpg manage 100 K 06 Jul 2017 - 13:02 DougMiller 123439 Modal
Topic revision: r2 - 14 Jul 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback